A model rocket is launched with an initial upward velocity of
60/ms. The rocket's height h (in meters) after t seconds is given by the following.
h=60t5t2
Find all values of t for which the rocket's height is 27 meters.

Answer by ikleyn(52778)

.
The correct equation is (and must be !) 


h(t) = -5t^2 + 60t.


To answer the question, you need to solve the equation  h(t) = 27,  which is the same as


-5t^2 + 60t = 27,


5t^2 - 60t + 27 = 0.


Solve it using the quadratic formula.


You will get two real positive roots.


One root is the time for  h(t) = 27 m  flying up;  the other root is the time for  h(t) = 27 m  falling down.






Plot h(t) = -5t^2 + 60t (red) and y = 27 (green)